# A complete proof of the geometric Bombieri–Lang conjecture for ramified covers of abelian varieties

### v1

@@ Abstract
We prove the geometric Bombieri–Lang conjecture for projective varieties which have finite maps to abelian varieties over function fields of characteristic 0. This generalizes the recent results of Xie–Yuan, which require either the hyperbolicity assumption or the non-isotriviality assumption. The proof is based on their strategy for constructing entire curves, but requires some new ingredients from algebraic geometry and Nevanlinna theory.
@@ 1 Introduction
The far-reaching Bombieri–Lang conjecture is a higher-dimensional generalization of the celebrated Mordell conjecture proved by Faltings [Fal1]. Beyond the Mordell conjecture, an important known case of the Bombieri–Lang conjecture is the case of subvarieties of abelian varieties by Faltings [Fal2, Fal3].
The geometric Bombieri–Lang conjecture is an analogue of the Bombieri–Lang conjecture over function fields. There are several versions of it. In the case of characteristic 0, [XY1, Conjecture 1.1] formulates a version that gives a very specific description of the structure of the rational points on a projective variety (over the function field). This version also implies most of other versions (see [XY1, §2.4] for more details). Thus we will use this version and focus on the case of characteristic 0 throughout this article.
The conjecture (or its suitable versions) is proved for curves by [Man, Gra], for smooth projective varieties with ample cotangent bundles by [Nog1, MD], for subvarieties of abelian varieties by [Ray, Par, Bui, Hru], and for constant hyperbolic varieties by [Nog2, Corollary 4.2].
In [XY1, XY2], Xie and Yuan gave a new idea for the general conjecture. Using this idea, they considered the case that $X$ has a finite map $f\colon X\to A$ to an abelian variety $A$ over $K$. Here $K$ is a finitely generated field over a field $k$ of characteristic 0. They proved in [XY1] the case when $X$ is hyperbolic, and in [XY2] the case when the $K/k$-trace $A^{(K/k)}$ of $A$ is 0. We refer to [Con] for more details on Chow’s trace of abelian varieties. It is worth mentioning that Parshin’s method in [Par] can also be used to prove the intersection of these two cases; see [BJ] for more details. Besides, [CDJLZ, Jav] also provides some results on the sparsity of rational points for ramified covers of abelian varieties.
The goal of this article is to remove both of the conditions above. Namely, we neither assume that $X$ is hyperbolic nor that $A$ is traceless. The approach is based on the strategy for constructing entire curves in [XY1, XY2], but requires new ingredients from algebraic geometry and Nevanlinna theory. We will state the main result in §1.1 and introduce the ideas in §1.2.
@@ 1.1 The main result
Now we state the main result. By a variety, we mean an integral separated scheme of finite type over a field. Recall that the algebraic special set ${\rm Sp}(X)$ [^1] In [XY1, XY2], they used ${\rm Sp_{alg}}(X)$ for the algebraic special set, we omit the subscript “alg” in this article. of a projective variety $X$ over a field $K$ is the Zariski closure in $X$ of the union of the images of all non-constant rational maps from abelian varieties over $\overline{K}$ to $X_{\overline{K}}$.
@@
Roughly speaking, the theorem says that all the rational points of $X$ outside the special locus come from the $k$-points of the $T_{i}$’s. Note that the set $\{Z_{1},\dots,Z_{r}\}$ contains all irreducible components of $Z$, but it may contain more elements to cover all rational points as in $(2)$; see [XY1, Example 2.12] for an example.
Theorem 1.1 confirms the geometric Bombieri–Lang conjecture formulated in [XY1, Conjecture 1.1] when $X$ has a finite map $f\colon X\to A$ to an abelian variety $A$. Furthermore, it generalizes [XY1, Theorem 1.2] and [XY2, Theorem 1.1] where one requires $X$ to be hyperbolic (i.e. ${\rm Sp}(X)=\emptyset$) and the other requires the $K/k$-trace of $A$ is 0. We refer to [Con] for more details on Chow’s trace of abelian varieties. Note that we do not assume that $f$ is surjective in the theorem. In particular, $X$ is allowed to be a closed subvariety of $A$, and in this case, versions of the geometric Bombieri–Lang conjecture were proved by [Fal3, Ray, Par, Bui, Hru].
We have the following corollary, which recovers [XY2, Theorem 1.1].
@@
Let $T_{i},Z_{i}$ be as in the statement of Theorem 1.1. By [XY1, Lemma 4.9], after a finite extension of $K$, the image of $T_{i}\times_{k}K\to Z_{i}\to X\to A$ is contained in some translation $(A^{(K/k)})_{K}+a_{i}\subseteq A$ where $a_{i}\in A(K)$ for all $i$. This proves the first statement. Now if $X$ is of general type and $A$ is not defined over $k$, then ${\rm Sp}(X)\subsetneqq X$ by [XY1, Corollary 4.3] (which is essentially due to Ueno and Kawamata) and $(A^{(K/k)})_{K}\subsetneqq A$, which implies that $X(K)$ is not Zariski dense in $X$. The third statement is obvious from the first one. ∎
To prove Theorem 1.1, we will first prove (in §4) the following variant, which generalizes [XY2, Corollary 1.2]. Recall that a normal projective $K$-variety $X$ is said to have a maximal Albanese dimension if the Albanese morphism $X_{\overline{K}}\to{\rm Alb}(X_{\overline{K}})$ (via a base point of $X(\overline{K})$) is generically finite onto its image. For convenience, we say a projective $K$-variety $X$ has a maximal Albanese dimension if the normalization of $X$ has.
@@
Let $K$ be a finitely generated field over a field $k$ of characteristic 0 such that $k$ is algebraically closed in $K$. Let $X$ be a projective variety of general type over $K$. Assume that $X$ has a maximal Albanese dimension. If $X(K)$ is Zariski dense in $X$, then the Albanese variety of (the normalization of) $X$ is defined over $k$ and $X$ is birational to $T_{K}$ for some projective variety $T$ over $k$. Moreover, the complement of the image of ${\rm Im}(T(k)\to T_{K}(K))$ in $X(K)$ ((via a birational map from $T_{K}$ to $X)$ is not Zariski dense in $X$.
If in addition the Albanese morphism of (the normalization of) $X$ is finite, then the normalization of $X$ is isomorphic to the base change $T_{K}=T\times_{k}K$ for a normal projective variety $T$ over $k$.
1.2 Ideas of the proof
Now we describe the idea to prove Theorem 1.1. The essential thing to prove is the following proposition.
Proposition 1.4.
@@ 1.2 Ideas of the proof
In fact, Proposition 1.4 above seems stronger than Proposition 3.1 in §3, but their proofs have no essential difference. The derivation of the main Theorem 1.1 from Proposition 3.1 will be given in §4, involving reduction steps similar to those in [XY1, XY2].
@@
In Proposition 1.4, if we take $T$ to be trivial and $f=(f_{0})_{K}$ arises as the base change of a morphism $f_{0}\colon X_{0}\to A_{0}$ of complex varieties, then $x_{n}$ can be seen as a holomorphic map $x_{n}\colon B\to X_{0}$ and the boundedness of the height of $\{x_{n}\}_{n\geq 1}$ is equivalent to the boundedness of the degree of $\{x_{n}(B)\}_{n\geq 1}$ in $X_{0}$ (with respect to some ample line bundle on $X_{0}$). This is known by [Yam2, Corollary 1(3)]. Therefore, Proposition 1.4 is a generalization of this result of Yamanoi.
Now we focus on the proof of Proposition 1.4. For convenience, we assume that $A$ has trivial $K/\mathbb{C}$-trace. Moreover, assume the compositions $X\to A\times_{K}T_{K}\to A$ and $X\to A\times_{K}T_{K}\to T_{K}$ are both surjective. The general situation is essentially the same. Let $f(x_{n})=(s_{n},t_{n})$ be the image of $x_{n}$ in $A(K)\times T_{K}(K)$. Then by our assumption, $\{s_{n}\}_{n\geq 1}$ is generic in $A$ and $\{t_{n}\}_{n\geq 1}$ is generic in $T$. By the condition in the proposition, we see that $t_{n}\in T(\mathbb{C})$ for all $n\geq 1$. By compactness, after taking a subsequence we may assume that $t_{n}$ tends to $t_{\infty}\in T(\mathbb{C})$ in the Euclidean topology.
@@
When $T$ is trivial, the image of the special set $Z={\rm Sp}(X)$ in $A$ is properly contained in $A$. As $\{x_{n}\}_{n\geq 1}$ is generic (in $X$), $\{s_{n}\}$ is also generic (in $f(X)$). After removing finitely many points, we can assume that $s_{n}\notin f(Z)$ for all $n$. Based on the unboundedness of the height of $\{s_{n}\}_{n\geq 1}$, in [XY2] Xie and Yuan constructed a linear entire curve in some fibre ${\mathcal{A}}_{b}\,(b\in B(\mathbb{C}))$ which is the limit of some re-parametrization of the sections $\{s_{n}\}_{n\geq 1}\colon B\to{\mathcal{A}}$ such that this entire curve is not contained in $f({\mathcal{Z}}_{b})$, where $\mathcal{Z}$} is the Zariski closure of $Z$ in $\mathcal{X}$}. Then they used Brody lemma to lift this entire curve to an entire curve on ${\mathcal{X}}_{b}$ which is not contained in ${\mathcal{Z}}_{b}$. On the other hand, [XY1] showed that the $b$ above can be chosen so that the analytic special set [^2] For a projective variety $X$ over $\mathbb{C}$, the analytic special set of $X$ is defined to be the Zariski closure of the union of the images of all nonconstant holomorphic maps $\phi\colon\mathbb{C}\to X$ in $X$. of ${\mathcal{X}}_{b}$ equals to ${\mathcal{Z}}_{b}$. This leads to a contradiction.
@@
However, when $T$ is not trivial, the situation gets more complicated. In this case, the image of the special set $Z={\rm Sp}(X)$ in $A\times_{K}T_{K}$ may contain $A\times\{t_{\infty}\}$. If this happens, any limit point of the sections $(s_{n},t_{n})\colon B\to{\mathcal{A}}\times_{\mathbb{C}}T(n\geq 1)$ is contained in ${\mathcal{A}}\times\{t_{\infty}\}\subseteq f({\mathcal{Z}})$. Thus the strategy in [XY2] is not sufficient.
To treat the general case, we will consider two complementary cases according to whether $f$ is surjective or not to deduce a contradiction.
The case when $f$ is not surjective
We first consider the case when $f$ is not surjective. In this case the method of Xie–Yuan still works after some efforts.
In fact, denote $Y=f(X)\subseteq A\times_{K}T_{K}$. Let $\mathcal{Y}$} be the Zariski closure of $Y$ in ${\mathcal{A}}\times_{\mathbb{C}}T$. Since $f$ is not surjective, $Y$ is a proper closed subset of $A\times_{K}T_{K}$ and $\mathcal{Y}$} is a proper closed subset of ${\mathcal{A}}\times_{\mathbb{C}}T$. If $\mathcal{Y}$} is flat over $T$, then ${\mathcal{Y}}_{t}$ is a proper closed subset of ${\mathcal{A}}\times\{t\}$ for all $t\in T$. Using [XY2, Theorem 3.3], we can construct a linear entire curve on ${\mathcal{A}}\times\{t_{\infty}\}$ (for some $t_{\infty}\in T(\mathbb{C})$) from $\{x_{n}\}_{n\geq 1}$ which is not contained in ${\mathcal{Y}}_{t_{\infty}}$. This leads to a contradiction because this entire curve is the limit of a re-parametrization of $\{(s_{n},t_{n})\colon B\to{\mathcal{Y}}\}_{n\geq 1}$, which must lie in $\mathcal{Y}$} as $\mathcal{Y}$} is closed in ${\mathcal{A}}\times_{\mathbb{C}}T$.
If $\mathcal{Y}$} is not flat over $T$, we can use a flattening theorem of Raynaud and Gruson (cf. [RG, Théorème 5.2.2]) to make it flat. This process does not change the truth of the proposition. This explains the proof when $f$ is not surjective. See §3.4 for more details.
The case when $f$ is surjective
Now we turn to the more subtle case when $f$ is surjective. In this situation Xie–Yuan’s method cannot directly deduce a contradiction.
@@
The main new ingredient in the proof of this case is a “tangency” observation. To illustrate the idea succinctly and intuitively, we temporarily assume that $X$ is smooth. In general we cannot make this assumption, but there is no big difference in this step. We will come back to the difficulty of singular case later.
@@
Now we assume Proposition 1.6. By Hurwitz’s formula, $K_{X}\sim f^{*}K_{(A\times_{K}T_{K})}+R$. Here $K_{X}$ is a canonical divisor of $X$, and similarly for $K_{(A\times_{K}T_{K})}$. Since the restriction of $f^{*}K_{(A\times_{K}T_{K})}$ to any fibre $X_{t}\,(t\in T)$ is trivial, we get that the restrictions of $K_{X}$ and $R$ to any fibre $X_{t}$ are linearly equivalent. Since $X$ is of general type, $K_{X}$ is big on $X$, and thus $R$ is a big divisor on the fibre $X_{t}$ for general $t\in T$. This positivity of $R$, combined with the fact that the height of $x_{n}$ tends to infinity as $n\to\infty$, allows us to “intuitively” imagine that $\mathcal{R}$} intersects sections $x_{n}\colon B\to{\mathcal{X}}$ at more and more points as $n\to\infty$. By Proposition 1.6, this is equivalent to the fact that ${\mathcal{D}}_{\rm red}$ is tangent to sections $(s_{n},t_{n})\colon B\to{\mathcal{A}}\times_{\mathbb{C}}T$ at more and more points as $n\to\infty$. Passing to the limit, we can imagine that the limit entire curve constructed in [XY2] is tangent to ${\mathcal{D}}_{\rm red}$ at “an excessive number of” points. As the limit entire curve is contained in the fibre ${\mathcal{A}}\times\{t_{\infty}\}$, it is tangent to $({\mathcal{D}}_{\rm red})_{t_{\infty}}$ at “an excessive number of” points.
@@
Now if $({\mathcal{D}}_{\rm red})_{t_{\infty}}$ is reduced and not equal to the whole fibre ${\mathcal{A}}\times\{t_{\infty}\}$, we get an entire curve on ${\mathcal{A}}\times\{t_{\infty}\}$ which is tangent to the reduced divisor $({\mathcal{D}}_{t_{\infty}})_{\rm red}$ on ${\mathcal{A}}\times\{t_{\infty}\}$ at “an excessive number of” points. Since $A$ has trivial trace, using [XY2, Theorem 3.3], we can construct this limit entire curve so that it is not contained in ${\mathcal{D}}_{b,t_{\infty}}$. Moreover, we can choose $b\in B(\mathbb{C})$ such that the fibre of the morphism $({\mathcal{D}}_{t_{\infty}})_{\rm red}\to B$ at $b$ is reduced. In this case we can deduce that the limit entire curve is tangent to the reduced divisor $({\mathcal{D}}_{b,t_{\infty}})_{\rm red}$ on ${\mathcal{A}}_{b}\times\{t_{\infty}\}$ at “an excessive number of” points. However, it is strange that a linear entire curve in a complex abelian variety is tangent at “an excessive number of” points to a reduced divisor which does not contain it. This leads to a contradiction.

### v2

# arXiv:2511.17010v2[math.NT] 02 Mar 2026

@@ Abstract
We prove the geometric Bombieri–Lang conjecture for projective varieties which have finite maps to abelian varieties over function fields of characteristic 0. This generalizes the recent results of Xie–Yuan, which require either the hyperbolicity assumption or the non-isotriviality assumption. The proof builds upon their strategy for constructing entire curves, yet hinges crucially on a new counting argument and draws substantially on tools from algebraic geometry and Nevanlinna theory to overcome various technical difficulties.
@@ 1 Introduction
The far-reaching Bombieri–Lang conjecture is a higher-dimensional generalization of the celebrated Mordell conjecture proved by Faltings [FAL83]. Beyond the Mordell conjecture, an important known case of the Bombieri–Lang conjecture is the case of subvarieties of abelian varieties by Faltings [FAL91, FAL94].
The geometric Bombieri–Lang conjecture is an analogue of the Bombieri–Lang conjecture over function fields. There are several versions of it. In the case of characteristic 0, [34, Conjecture 1.1] formulates a version that gives a very specific description of the structure of the rational points on a projective variety (over the function field). This version also implies most of other versions (see [34, §2.4] for more details). Thus we will use this version and focus on the case of characteristic 0 throughout this article.
The conjecture (or its suitable versions) is proved for curves by [MAN63, GRA65], for smooth projective varieties with ample cotangent bundles by [NOG81, MAR84, GR18], for subvarieties of abelian varieties by [RAY83, PAR90, BUI92, HRU96], and for constant hyperbolic varieties by [NOG92, Corollary 4.2]. See [BOG77] (or [GAS02, Theorem 2.3]), [20, Theorem E, Theorem 6.3] for further results on constant varieties.
Recently, Xie and Yuan considered the case that $X$ has a finite map $f\colon X\to A$ to an abelian variety $A$ over $K$. Here $K$ is a finitely generated field over a field $k$ of characteristic 0. They proved in [34] the case when $X$ is hyperbolic, and in [35] the case when the $K/k$-trace $A^{(K/k)}$ of $A$ is 0. We refer to [CON06] for more details on Chow’s trace of abelian varieties. It is worth mentioning that Parshin’s method in [PAR90] can also be used to prove the intersection of these two cases; see [BJ26] for more details. Besides, [CDJ+22, 21] also provides some results on the sparsity of rational points for ramified covers of abelian varieties.
The goal of this article is to remove both of the conditions above. Namely, we neither assume that $X$ is hyperbolic nor that $A$ is traceless. Unlike most problems over function fields, a non-trivial trace creates essential difficulty here; see the discussion in §1.2. Our approach starts from the strategy for constructing entire curves in [35], but hinges on a new counting argument, which leads to several new technical difficulties. To overcome these difficulties, we require extensive use of tools from algebraic geometry and Nevanlinna theory. We will state the main result in §1.1 and introduce the ideas in §1.2, §1.3 and §1.4. Specifically, §1.2 introduces some basic reductions and setups, §1.3 then presents the key counting argument. Finally §1.4 addresses three technical difficulties and their resolutions.
@@ 1.1 The main result
Now we state the main result. By a variety, we mean an integral separated scheme of finite type over a field. Recall that the algebraic special set ${\rm Sp}(X)$ [^1] In [34, 35], they used ${\rm Sp_{alg}}(X)$ for the algebraic special set, we omit the subscript “alg” in this article. of a projective variety $X$ over a field $K$ is the Zariski closure in $X$ of the union of the images of all non-constant rational maps from abelian varieties over $\overline{K}$ to $X_{\overline{K}}$.
@@
Roughly speaking, the theorem says that all the rational points of $X$ outside the special locus come from the $k$-points of the $T_{i}$’s. Note that the set $\{Z_{1},\dots,Z_{r}\}$ contains all irreducible components of $Z$, but it may contain more elements to cover all rational points as in $(2)$; see [34, Example 2.12] for an example.
Theorem 1.1 confirms the geometric Bombieri–Lang conjecture formulated in [34, Conjecture 1.1] when $X$ has a finite map $f\colon X\to A$ to an abelian variety $A$. Furthermore, it generalizes [34, Theorem 1.2] and [35, Theorem 1.1] where one requires $X$ to be hyperbolic (i.e. ${\rm Sp}(X)=\emptyset$) and the other requires the $K/k$-trace of $A$ is 0. We refer to [CON06] for more details on Chow’s trace of abelian varieties. Note that we do not assume that $f$ is surjective in the theorem. In particular, $X$ is allowed to be a closed subvariety of $A$, and in this case, versions of the geometric Bombieri–Lang conjecture were proved by [FAL94, RAY83, PAR90, BUI92, HRU96].
We have the following corollary, which recovers [35, Theorem 1.1].
@@
Let $T_{i},Z_{i}$ be as in the statement of Theorem 1.1. By [34, Lemma 4.9], after a finite extension of $K$, the image of $T_{i}\times_{k}K\to Z_{i}\to X\to A$ is contained in some translation $(A^{(K/k)})_{K}+a_{i}\subseteq A$ where $a_{i}\in A(K)$ for all $i$. This proves the first statement. Now if $X$ is of general type and $A$ is not defined over $k$, then ${\rm Sp}(X)\subsetneqq X$ by [34, Corollary 4.3] (which is essentially due to Ueno and Kawamata) and $(A^{(K/k)})_{K}\subsetneqq A$, which implies that $X(K)$ is not Zariski dense in $X$. The third statement is obvious from the first one. ∎
To prove Theorem 1.1, we will first prove (in §4) the following variant, which generalizes [35, Corollary 1.2]. Recall that a normal projective $K$-variety $X$ is said to have a maximal Albanese dimension if the Albanese morphism $X_{\overline{K}}\to{\rm Alb}(X_{\overline{K}})$ (via a base point of $X(\overline{K})$) is generically finite onto its image. For convenience, we say a projective $K$-variety $X$ has a maximal Albanese dimension if the normalization of $X$ has.
@@
Let $K$ be a finitely generated field over a field $k$ of characteristic 0 such that $k$ is algebraically closed in $K$. Let $X$ be a projective variety of general type over $K$. Assume that $X$ has a maximal Albanese dimension. If $X(K)$ is Zariski dense in $X$, then the Albanese variety of ((the normalization of) $X$ is defined over $k$ and $X$ is birational to $T_{K}$ for some projective variety $T$ over $k$. Moreover, the complement of the image of ${\rm Im}(T(k)\to T_{K}(K))$ in $X(K)$ ((via a birational map from $T_{K}$ to $X)$ is not Zariski dense in $X$.
If in addition the Albanese morphism of ((the normalization of) $X$ is finite, then the normalization of $X$ is isomorphic to the base change $T_{K}=T\times_{k}K$ for a normal projective variety $T$ over $k$.
1.2 Basic setups
Now we present the reductions and setups for the proof of Theorem 1.1. The essential thing to prove is the following proposition, which is the goal of §3.
Proposition 1.4 (Proposition 3.1).
@@ 1.2 Basic setups
The derivation of the main Theorem 1.1 from Proposition 1.4 will be given in §4, involving reduction steps similar to those in [34].
@@
(1) We explain that $T$ in the proposition plays the role of the “constant part”. Even if we are only concerned with finite covers of the abelian variety $A$ (i.e. the case where $T$ is trivial), when the $K/\mathbb{C}$-trace of $A$ is nontrivial, we sometimes transform the problem into the situation where $T$ is nontrivial. In this transformation, $T_{K}$ becomes a constant quotient abelian variety of $A$; see §3.3, “Basic setups” for more details.
(2) In Proposition 1.4, if we take $T$ to be trivial and $f=(f_{0})_{K}$ arises as the base change of a morphism $f_{0}\colon X_{0}\to A_{0}$ of complex varieties, then $x_{n}$ can be seen as a holomorphic map $x_{n}\colon B\to X_{0}$ and the boundedness of the height of $\{x_{n}\}_{n\geq 1}$ is equivalent to the boundedness of the degree of $\{x_{n}(B)\}_{n\geq 1}$ in $X_{0}$ (with respect to some ample line bundle on $X_{0}$). This is known by [YAM15a, Corollary 1(3)]. Therefore, Proposition 1.4 is a generalization of this result of Yamanoi.
Now we focus on the proof of Proposition 1.4. For convenience, we assume that $f$ is surjective and $A$ has trivial $K/\mathbb{C}$-trace. This actually captures the most essential case. Let $f(x_{n})=(s_{n},t_{n})$ be the image of $x_{n}$ in $A(K)\times T_{K}(K)$. Then by our assumption, $\{s_{n}\}_{n\geq 1}$ is generic in $A$ and $\{t_{n}\}_{n\geq 1}$ is generic in $T$. By the condition in the proposition, we see that $t_{n}\in T(\mathbb{C})$ for all $n\geq 1$. By compactness, after taking a subsequence we may assume that $t_{n}$ tends to $t_{\infty}\in T(\mathbb{C})$ in the Euclidean topology.
@@
When $T$ is trivial, the image of the special set $Z={\rm Sp}(X)$ in $A$ is properly contained in $A$. As $\{x_{n}\}_{n\geq 1}$ is generic (in $X$), $\{s_{n}\}$ is also generic (in $f(X)$). After removing finitely many points, we can assume that $s_{n}\notin f(Z)$ for all $n$. Based on the unboundedness of the height of $\{s_{n}\}_{n\geq 1}$, in [35] Xie and Yuan constructed a linear entire curve in some fibre ${\mathcal{A}}_{b}\,(b\in B(\mathbb{C}))$ which is the limit of some re-parametrization of the sections $\{s_{n}\}_{n\geq 1}\colon B\to{\mathcal{A}}$ such that this entire curve is not contained in $f({\mathcal{Z}}_{b})$, where $\mathcal{Z}$} is the Zariski closure of $Z$ in $\mathcal{X}$}. Then they used Brody lemma to lift this entire curve to an entire curve on ${\mathcal{X}}_{b}$ which is not contained in ${\mathcal{Z}}_{b}$. On the other hand, [34] showed that the $b$ above can be chosen so that the analytic special set [^2] For a projective variety $X$ over $\mathbb{C}$, the analytic special set of $X$ is defined to be the Zariski closure of the union of the images of all nonconstant holomorphic maps $\phi\colon\mathbb{C}\to X$ in $X$. of ${\mathcal{X}}_{b}$ equals to ${\mathcal{Z}}_{b}$. This leads to a contradiction.
@@
However, when $T$ is not trivial, the situation gets more complicated. In this case, the image of the special set $Z={\rm Sp}(X)$ in $A\times_{K}T_{K}$ may contain $A\times\{t_{\infty}\}$. If this happens, any limit point of the sections $(s_{n},t_{n})\colon B\to{\mathcal{A}}\times_{\mathbb{C}}T(n\geq 1)$ is contained in ${\mathcal{A}}\times\{t_{\infty}\}\subseteq f({\mathcal{Z}})$. Thus the strategy in [35] is not sufficient.
New idea
To treat the general case, we will employ a rather different approach—a counting argument. We still use the construction of linear entire curves and control their positions. But instead of studying the positions of the entire curves relative to ${\rm Sp}(X)$, we focus on the ramification divisor $R$ of $f$. We will derive a contradiction by estimating the intersection numbers of the sections $\{x_{n}\}_{n\geq 1}$ with the Zariski closure $\mathcal{R}$} of $R$ in $\mathcal{X}$}. Notably, this approach also yields a new proof for [35, Theorem 1.1].
Remark 1.6.
In fact, we will not use anything about the special set ${\rm Sp}(X)$ in the proof of Proposition 1.4 (and also Proposition 3.1). The only thing we will use for it is in the reduction steps in §4, especially the Green–Griffiths–Lang conjecture for finite covers of abelian varieties, which is purely algebraic geometry. Therefore, although both [35] and the present article use the construction of linear entire curves and require controlling their positions, their underlying philosophies are quite different.
1.3 Key idea: a counting argument
Now we begin to explain the crucial counting argument. We will first present the geometric intuition and then describe the method for its rigorous formulation.
Tangency
@@ 1.3 Key idea: a counting argument
The idea originates from a “tangency” observation. To illustrate it succinctly and intuitively, we temporarily assume that $X$ is smooth. In general we cannot make this assumption, but there is no big difference in this step. We will come back to the difficulty of singular case in §1.4.
@@
Now we assume Proposition 1.7. By Hurwitz’s formula, $K_{X}\sim f^{*}K_{(A\times_{K}T_{K})}+R$. Here $K_{X}$ is a canonical divisor of $X$, and similarly for $K_{(A\times_{K}T_{K})}$. Since the restriction of $f^{*}K_{(A\times_{K}T_{K})}$ to any fibre $X_{t}\,(t\in T)$ is trivial, we get that the restrictions of $K_{X}$ and $R$ to any fibre $X_{t}$ are linearly equivalent. Since $X$ is of general type, $K_{X}$ is big on $X$, and thus $R$ is a big divisor on the fibre $X_{t}$ for general $t\in T$.
This positivity of $R$, combined with the fact that the height of $x_{n}$ tends to infinity as $n\to\infty$, allows us to “intuitively” imagine that $\mathcal{R}$} intersects sections $x_{n}\colon B\to{\mathcal{X}}$ at more and more points as $n\to\infty$. By Proposition 1.7, this implies that ${\mathcal{D}}_{\rm red}$ is tangent to sections $(s_{n},t_{n})\colon B\to{\mathcal{A}}\times_{\mathbb{C}}T$ at more and more points as $n\to\infty$. Passing to the limit, we can imagine that the limit entire curve constructed in [35] is tangent to ${\mathcal{D}}_{\rm red}$ at “an excessive number of” points. As the limit entire curve is contained in the fibre ${\mathcal{A}}\times\{t_{\infty}\}$, it is tangent to $({\mathcal{D}}_{\rm red})_{t_{\infty}}$ at “an excessive number of” points.
@@
Now if $({\mathcal{D}}_{\rm red})_{t_{\infty}}$ is reduced and not equal to the whole fibre ${\mathcal{A}}\times\{t_{\infty}\}$, we get an entire curve on ${\mathcal{A}}\times\{t_{\infty}\}$ which is tangent to the reduced divisor $({\mathcal{D}}_{t_{\infty}})_{\rm red}$ on ${\mathcal{A}}\times\{t_{\infty}\}$ at “an excessive number of” points. Since $A$ has trivial trace, using [35, Theorem 3.3], we can construct this limit entire curve so that it is not contained in ${\mathcal{D}}_{b,t_{\infty}}$. Moreover, we can choose $b\in B(\mathbb{C})$ such that the fibre of the morphism $({\mathcal{D}}_{t_{\infty}})_{\rm red}\to B$ at $b$ is reduced. In this case we can deduce that the limit entire curve is tangent to the reduced divisor $({\mathcal{D}}_{b,t_{\infty}})_{\rm red}$ on ${\mathcal{A}}_{b}\times\{t_{\infty}\}$ at “an excessive number of” points. However, it is strange that a linear entire curve in a complex abelian variety is tangent at “an excessive number of” points to a reduced divisor which does not contain it. This leads to a contradiction.
In general, we have to handle the case when $({\mathcal{D}}_{\rm red})_{t_{\infty}}$ is non-reduced or it is the whole fibre ${\mathcal{A}}\times\{t_{\infty}\}$. This leads to the other two technical difficulties and will be treated in the next subsection.
Rigorous realization via Nevanlinna theory
The above is the rough idea, but it is far from rigorous. For example, the meaning of “an excessive number of” points is not clear. We will use Nevanlinna theory (see §3.1) to make the proof rigorous.
Specifically, we will employ the counting function in Nevanlinna theory to quantify the number of intersections between $x_{n}(B)$ and $\mathcal{R}$}, as well as the number of tangent points between $(s_{n},t_{n})(B)$ and ${\mathcal{D}}_{\rm red}$. Eventually, we will show two things: on the one hand, the positivity of $R$ implies that the tangent points between $(s_{n},t_{n})(B)$ and ${\mathcal{D}}_{\rm red}$ account for a strictly positive proportion of all intersection points (in the asymptotic sense; see Lemma 3.18); on the other hand, by Yamanoi’s Second Main Theorem (see Theorem 3.6(1)) and some limit arguments (Lemma 1.9 below and Lemma 3.5), we will prove that this proportion of tangent points must be zero (see Lemma 3.16), thereby derive a contradiction.
Note that $\{(s_{n},t_{n})\}_{n\geq 1}$ are merely sections, not entire curves. The limit entire curve is obtained by re-parametrizing each $(s_{n},t_{n})$ by open discs of increasing radii and then passing to limit (see Theorem 2.2). Therefore, unlike the classical Nevanlinna theory, we have to consider the counting functions (as well as proximity and characteristic functions) for holomorphic maps from an open disc to a complex variety, rather than only from the whole complex plane. Furthermore, we need to vary the holomorphic maps and study the limit behaviour of these three functions in Nevanlinna theory (see Lemma 3.5).
Remark 1.8.
In practice, it is more convenient to consider asymptotic behaviour rather than limiting behaviour. Specifically, in the two opposite estimates above, one is easier to handle for each individual section $(s_{n},t_{n})$ (the asymptotic behaviour), while the other fits naturally with the limit entire curve (the limiting behaviour). Hence, in order to get a contradiction, we have two possible routes: either pass from the asymptotic behaviour to the limiting behaviour, or conversely.
Classically, it is more natural to transform both estimates into bounds for the limit entire curve, since more tools are available for entire curves. In our setting, however, the second route—from the limit to the asymptotic—proves to be easier. The essential reason is that the limit entire curve exists only on ${\mathcal{A}}\times_{\mathbb{C}}T$ (it can also be lifted to $\mathcal{X}$}, but its behaviour becomes bad when the fibre ${\mathcal{X}}_{t_{\infty}}$ lies in the singular locus of $\mathcal{X}$}), whereas the tangency comes from intersections of $x_{n}$ and $\mathcal{R}$}, which live on $\mathcal{X}$}. Hence we need to use the limit entire curve as a bridge to obtain information about the asymptotic behaviour.
1.4 Three technical difficulties
As we have mentioned in the last subsection, there are three difficulties in realizing the counting argument above. We now explain each in turn.
Non-reducedness of $({\mathcal{D}}_{\rm red})_{t_{\infty}}$